# Representation-protected topology of spin-singlet $s$-wave superconductors

### v1

@@ Introduction.
The K-theoretical classification based on stable equivalence fails to capture some topological phases. The topological phases that fall outside the K-theoretical classification are called unstable topological phases and can become topologically nontrivial only if certain conditions are met regarding the number of occupied (or total) bands or band representations. The search for unstable topological phases is vigorously underway and has revealed a variety of unstable topological phases such as fragile [34, 35] and delicate [36, 37] topological insulators. For example, the Hopf insulators [27, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] are in a delicate topological phase and realized only when the total number of bands is limited to two, while the Euler insulators [48, 49, 50, 51, 52] are fragile topological insulators that are realized only when the number of valence bands is limited to two.
@@
Furthremore, extending the definition of the $\mathbb{Z}_{2}$ index to three-dimensional (3D) superconductors, we obtain a relationship between the Fermi surface topology of a 3D $s_{\pm}$-wave superconductor and a 3D winding number ($w_{3d}\in 2\mathbb{Z}$), a stable topological invariant in class CI [27]. This gives the following Fermi surface criterion of the 3D topological phase: $w_{3d}\neq 0$ if the two Fermi surfaces of the orbital doublets, each enclosing one of the $C_{n}T$ invariant momenta, have $s_{\pm}$-wave pair potentials. Since the 3D winding number is defined independently of the $C_{n}$ symmetry and band representations, the topological phases with $w_{3d}\neq 0$ are robust against perturbations which break the prerequisites (i) and (ii) as long as the bulk energy gap and the defining symmetries of class CI are maintained. We construct a model of a 3D representation-protected $s_{\pm}$-wave topological superconductor that satisfies the Fermi surface criterion, and demonstrate the existence and robustness of gapless surface states with quadratic dispersion.
@@
$H(\bm{k}_{\parallel})=\begin{pmatrix}0&A(\bm{k}_{\parallel})\\ A^{\dagger}(\bm{k}_{\parallel})&0\end{pmatrix},$ (4)
where $\bm{k}_{\parallel}=(k_{x},k_{y})$ and $A(\bm{k}_{\parallel})\equiv\mathcal{E}(\bm{k}_{\parallel})-i\Delta(\bm{k}_{\parallel})$. The $C_{4,1}T$ operator acting on the off-diagonal component $A(\bm{k}_{\parallel})$ satisfies the relation
@@
The expression (6) is gauge invariant and only takes the values ±$\pm 1$ [32]. Thus, a spin-singlet superconducting state with $\nu_{4}=1$ cannot be continuously transformed to any spin-single superconductor with $\nu_{4}=0$ without closing the superconducting gap. Since Eq. (6) is defined solely in terms of the orbital doublets, hybridization with an $s$-orbital band makes Eq. (6) ill-defined. The topological superconductor of orbital doublets is a superconducting analog of the class AI representation-proteccted topological insulator proposed by Fu [57].
@@
Let us assume that 3D superconductors of our interest have $C_{4}$ rotation symmetry along the $z$ axis. The 3D systems have two $C_{4}T$ symmetric planes $k_{z}=0$ and $\pi$, on which we can define the $\mathbb{Z}_{2}$ topological invariants independently, labeled by $\nu_{4}(0)$ and $\nu_{4}(\pi)$. The $C_{4}T$ invariant momenta are shown in Fig. 1 (b). The $\nu_{4}(0)$ and $\nu_{4}(\pi)$ are defined on the $\Gamma$-M line and the Z-A line, respectively. Then, 3D topological phases are characterized by the difference $\tilde{\nu}_{4}=\nu_{4}(\pi)-\nu_{4}(0)\mod 2$. We call the phase with $\tilde{\nu}_{4}=1$ a strong topological phase. The strong topological phase has a gapless topological surface state on the (001) surface that is invariant under the $C_{4}$ symmetry. On the other hand, when $\tilde{\nu}_{4}=0$, we have either a trivial phase or a weak topological phase, i.e., a stack of 2D topological phases.
Furthermore, there are stable topological phases characterized by the 3D winding number $w_{3d}$ in class CI. We find that $\tilde{\nu}_{4}$ and $w_{3d}$ are related by
@@
In the weak-coupling limit, the 3D winding number is written as $w_{\rm 3d}=\frac{1}{2}\sum_{s}\operatorname{sgn}(\delta_{k_{F},s})C_{s}$ [32], where $C_{s}$ is the Chern number of the $s$th Fermi surface, and the summation is taken over all the Fermi surfaces. The derivation of Eq. (8) is given in SM [60]. Equation (8) is one of our main results, bridging the representation-protected $\mathbb{Z}_{2}$ topology and the stable topology. Note that the formula is meaningful only in the absence of inversion symmetry, since the inversion symmetry leads to $w_{\rm 3d}=0$ [61].
@@
We now turn to $C_{6}$ symmetric cases. The 2D and 3D Brillouin zones are shown in Fig 1 (c) and (d), where $C_{6}T$ invariant momenta are denoted by $\Gamma$, Z, K, and H. After repeating the same procedure as the $C_{4}$ symmetric case, we obtain the weak $\mathbb{Z}_{2}$ topological invariant as
@@ Discussion.
We have shown that the representation-protected $\mathbb{Z}_{2}$ topological invariants [Eqs. (6) and (10)] are defined for fully-gapped superconductors with SU(2) spin rotation, $C_{n}$ rotation ($n=4,6$), and time-reversal symmetries, when the Fermi surfaces consist only of the orbital doublets with the angular momentum $\ell=1$ for $C_{4}$ and $\ell=1,2$ for $C_{6}$. We have established the relation between the $\mathbb{Z}_{2}$ index and the 3D winding number via the Fermi-surface formula (7) and (8) in the weak-coupling limit. We have presented a model of 3D topological $s_{\pm}$ superconductors, which has a gapless surface mode with quadratic dispersion. Finally, we have proposed pressurized KFe2As2 as a potential candidate of a representation-protected topological superconductor in class CI.

### v2

@@ Introduction.
The K-theoretical classification based on stable equivalence fails to capture some topological phases. The topological phases that fall outside the K-theoretical classification are called unstable topological phases and can become topologically nontrivial only if certain conditions are met regarding the number of occupied (or total) bands or band representations. The search for unstable topological phases is vigorously underway and has revealed a variety of unstable topological phases such as fragile [34, 35] and delicate [36, 37] topological insulators. For example, the Hopf insulators [27, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] are in a delicate topological phase and realized only when the total number of bands is limited to two, while the Euler insulators [50, 51, 52, 53, 54, 55] are fragile topological insulators that are realized only when the number of valence bands is limited to two.
@@
Furthermore, extending the definition of the $\mathbb{Z}_{2}$ index to three-dimensional (3D) superconductors, we obtain a relationship between the Fermi surface topology of a 3D $s_{\pm}$-wave superconductor and a 3D winding number ($w_{3d}\in 2\mathbb{Z}$; see Eq. (8)), a stable topological invariant in class CI [27]. This gives the following Fermi surface criterion of the 3D topological phase: $w_{3d}\neq 0$ if the two Fermi surfaces of the orbital doublets, each enclosing one of the $C_{n}T$ invariant momenta, have $s_{\pm}$-wave pair potentials. Since the 3D winding number is defined independently of the $C_{n}$ symmetry and band representations, the topological phases with $w_{3d}\neq 0$ are robust against perturbations which break the prerequisites (i) and (ii) as long as the bulk energy gap and the defining symmetries of class CI are maintained. We construct a model of a 3D representation-protected $s_{\pm}$-wave topological superconductor that satisfies the Fermi surface criterion, and demonstrate the existence and robustness of gapless surface states with quadratic dispersion.
@@
$H(\bm{k}_{\parallel})=\begin{pmatrix}0&A(\bm{k}_{\parallel})\\ A^{\dagger}(\bm{k}_{\parallel})&0\end{pmatrix}$ (4)
in the basis of $(c_{\bm{k},\uparrow,\sigma}-ic^{\dagger}_{-\bm{k},\downarrow,\sigma},c_{\bm{k},\uparrow,\sigma}+ic^{\dagger}_{-\bm{k},\downarrow,\sigma})^{t}/\sqrt{2}$, where $\bm{k}_{\parallel}=(k_{x},k_{y})$ and $A(\bm{k}_{\parallel})\equiv\mathcal{E}(\bm{k}_{\parallel})-i\Delta(\bm{k}_{\parallel})$. The $C_{4,1}T$ operator acting on the off-diagonal component $A(\bm{k}_{\parallel})$ satisfies the relation
@@
The expression (6) is gauge invariant and only takes the values ±$\pm 1$ [32, 63]. Thus, a spin-singlet superconducting state with $\nu_{4}=1$ cannot be continuously transformed to any spin-singlet superconductor with $\nu_{4}=0$ without closing the superconducting gap. Since Eq. (6) is defined solely in terms of the orbital doublets, hybridization with an $s$-orbital band makes Eq. (6) ill-defined. The topological superconductor of orbital doublets is a superconducting analog of the class AI representation-protected topological insulator proposed by Fu [60].
@@
Let us assume that 3D superconductors of our interest have $C_{4}$ rotation symmetry along the $z$ axis. The 3D systems have two $C_{4}T$ symmetric planes $k_{z}=0$ and $\pi$, on which we can define the $\mathbb{Z}_{2}$ topological invariants independently, labeled by $\nu_{4}(0)$ and $\nu_{4}(\pi)$. The $C_{4}T$ invariant momenta are shown in Fig. 1 (b). The $\nu_{4}(0)$ and $\nu_{4}(\pi)$ are defined on the $\Gamma$-M line and the Z-A line, respectively. Then, 3D topological phases are characterized by the difference $\tilde{\nu}_{4}=\nu_{4}(\pi)-\nu_{4}(0)\mod 2$. We call the phase with $\tilde{\nu}_{4}=1$ a strong topological phase. The strong topological phase has a gapless topological surface mode on the (001) surface that is invariant under the $C_{4}$ symmetry. On the other hand, when $\tilde{\nu}_{4}=0$, we have either a trivial phase or a weak topological phase, i.e., a stack of 2D topological phases.
Furthermore, there are stable topological phases characterized by the 3D winding number $w_{3d}$ in class CI. The 3D winding number is defined as follows [27]. The BdG Hamiltonian of a 3D class CI superconductor can be transformed to the off-diagonal matrix form as in Eq. (4). Then, $A(\bm{k})$ is rewritten by singular value decomposition as $A(\bm{k})=U(\bm{k})D(\bm{k})V^{\dagger}(\bm{k})$ with the unitary matrices $U(\bm{k})$ and $V(\bm{k})$, where $\bm{k}=(k_{x},k_{y},k_{z})$. Since $D(\bm{k})$ is positive definite, we can deform it to the identity matrix without closing the energy gap. Therefore, we can replace $A(\bm{k})$ with a unitary matrix $Q(\bm{k})=U(\bm{k})V^{\dagger}(\bm{k})$ and define the 3D winding number
$w_{\rm 3d}=$ $\int\frac{d^{3}k}{24\pi^{2}}\sum_{\alpha,\beta,\gamma}\epsilon^{\alpha\beta\gamma}\operatorname{Tr}\Big{\{}\left[Q^{\dagger}(\bm{k})\partial_{k_{\alpha}}Q(\bm{k})\right]$
×$\times\left[Q^{\dagger}(\bm{k})\partial_{k_{\beta}}Q(\bm{k})\right]\!\left[Q^{\dagger}(\bm{k})\partial_{k_{\gamma}}Q(\bm{k})\right]\Big{\}},$ (8)
where $\alpha,\beta,\gamma=x,y,z$ and $\epsilon^{\alpha\beta\gamma}$ is the Levi-Civita symbol. The winding number $w_{\rm 3d}$ takes even-integer values in class CI. In the weak-coupling limit, it is reduced to the form $w_{\rm 3d}=\frac{1}{2}\sum_{s}\operatorname{sgn}(\delta_{k_{F},s})C_{s}$ [32], where $C_{s}$ is the Chern number of the $s$th Fermi surface, and the summation is taken over all the Fermi surfaces. Using this Fermi surface formula, we find that $\tilde{\nu}_{4}$ and $w_{\rm 3d}$ are related by
@@
The derivation of Eq. (9) is given in SM [63]. Equation (9) is one of our main results, bridging the representation-protected $\mathbb{Z}_{2}$ topology and the stable topology. Note that the formula is meaningful only in the absence of inversion symmetry, since the inversion symmetry leads to $w_{\rm 3d}=0$ [64].
@@
We now turn to $C_{6}$ symmetric cases. The 2D and 3D Brillouin zones are shown in Fig. 1 (c) and (d), where $C_{6}T$ invariant momenta are denoted by $\Gamma$, Z, K, and H. After repeating the same procedure as the $C_{4}$ symmetric case, we obtain the weak $\mathbb{Z}_{2}$ topological invariant as
@@ Discussion.
We expect from the Fermi surface criterion that materials with only electron pockets, such as KFe2Se2 [84, 85, 86], monolayer FeSe [87, 88, 89, 90, 91], and LiFe(OH)FeSe [92, 93], should also be potential candidates. Investigating representation-protected topological phases in other FeSCs is also an interesting future work.
@@
We have shown that the representation-protected $\mathbb{Z}_{2}$ topological invariants [Eqs. (6) and (11)] are defined for fully-gapped superconductors with SU(2) spin rotation, $C_{n}$ rotation ($n=4,6$), and time-reversal symmetries, when the Fermi surfaces consist only of the orbital doublets with the angular momentum $\ell=1$ for $C_{4}$ and $\ell=1,2$ for $C_{6}$. We have established the relation between the $\mathbb{Z}_{2}$ index and the 3D winding number via the Fermi-surface formula (7) and (9) in the weak-coupling limit. We have presented a model of 3D topological $s_{\pm}$ superconductors, which has a gapless surface mode with quadratic dispersion. Such gapless surface mode would be observed through surface-sensitive measurements, such as tunneling spectroscopy [94, 95], the Josephson effect [96, 97], and penetration depth measurement [98]. Finally, we have proposed pressurized KFe2As2 as a potential candidate of a representation-protected topological superconductor in class CI.